OpenStudy (anonymous):

how do you factor -x2y2+x4+9y2-9x2

OpenStudy (anonymous):

Delta_x = 144 y^2 (y^4-18 y^2+81)^2

OpenStudy (turingtest):

\[-x^2y^2+x^4+9y^2-9x^2=(x^2-y^2)(x^2-9)\]

OpenStudy (anonymous):

\[-x^2 y^2+x^4+9 y^2-9 x^2\text{=}(x-3) (x+3) (x-y) (x+y) \]